# MorphAny3D: Unleashing the Power of Structured Latent in 3D Morphing

### v1

@@ 4.2 Evaluation
3DInterp [densematcher] 409.14 2.55 0.0006 1.00 0.61
SLATInterp [densematcher] 348.31 6.53 0.0010 0.00 1.43
DiffMorpher [diffmorpher] 208.08 6.65 0.0021 5.00 0.82
FreeMorph [freemorph] 164.68 5.89 0.0027 11.00 3.27
@@ 4.5 Generalization Ability
We apply MorphAny3D to other SLAT-based models—Hi3DGen [hi3dgen] and Text-to-3D Trellis [trellis]. As shown in Fig. 10, it consistently produces smooth, high-quality morphs, confirming its versatility.
@@ 5 Conclusion
Limitations and Future Work. MorphAny3D inherits Trellis’s limitations and may exhibit artifacts on extremely fine structures (see red boxes in Fig. 11). Future work could adopt stronger 3D generative backbones [triposf] or enhance SLAT representations to capture finer geometric details.

### v2

@@ 4.2 Evaluation
3DInterp [94] 409.14 2.55 0.0006 1.00 0.61
SLATInterp [94] 348.31 6.53 0.0010 0.00 1.43
DiffMorpher [90] 208.08 6.65 0.0021 5.00 0.82
FreeMorph [4] 164.68 5.89 0.0027 11.00 3.27
@@ 4.5 Generalization Ability
We apply MorphAny3D to other SLAT-based models—Hi3DGen [86] and Text-to-3D Trellis [81]. As shown in Fig. 10, it consistently produces smooth, high-quality morphs, confirming its versatility.
@@ 5 Conclusion
Limitations and Future Work. MorphAny3D inherits Trellis’s limitations and may exhibit artifacts on extremely fine structures (see red boxes in Fig. 11). Future work could adopt stronger 3D generative backbones [28] or enhance SLAT representations to capture finer geometric details.